# Fast LISA likelihood approximations by downsampling

### v1

@@ 3.1.2 Residuals and metrics
There are a number of ways to go about expressing the downsampled residual in terms of the fully sampled residual. First, consider the $N_{\mathrm{s}}\times N_{\mathrm{f}}$ matrix
@@
for all $\mathbf{v}$, where $\bm{\Gamma}^{-1}_{\mathrm{s}}$ and $\bm{\Gamma}^{-1}_{\mathrm{f}}$ are inner products on the subspace and full data space respectively. An exact solution for $\bm{\Gamma}^{-1}_{\mathrm{s}}$ would necessarily be expressed in terms of $\mathbf{v}$. Our method will be to make an initial guess of a metric on the subspace ${\bm{\Gamma}^{\prime}}_{\mathrm{s}}^{-1}$, and define the final metric $\bm{\Gamma}^{-1}_{\mathrm{s}}=\bm{\Gamma}^{-1}_{\mathrm{s}}(\bm{\Gamma}^{\prime}_{\mathrm{s}})$ in terms of the initial guess in some appropriate fashion, as shall be described in the following sections. With ${\bm{\Gamma}^{\prime}}_{\mathrm{s}}^{-1}$ an an $N_{\mathrm{s}}\times N_{\mathrm{s}}$ matrix, we can project the full data space metric down to the subspace using $\mathbf{D}$ via
@@
3.1.3 Time domain sample correlations
@@
The above example shows how correlations cause the information in the time-domain signal to which the detector is sensitive to become mixed between neighbouring samples. In order to prevent this from impinging upon the Bayesian inference after downsampling, we must whiten, or in some other way decorrelate the data samples (for example, frequency representations from fast Fourier transform (FFT) s often consist of uncorrelated samples). Thus in fact the preceding prescription for downsampling given in Section 3.1.2 should only be carried out on uncorrelated data: since frequency domain data is naturally uncorrelated, the task would be straightforward and safe (at least from the point of view of sample correlations) to downsample frequency samples. In the time domain, we must first whiten residuals before downsampling. This simply means setting
@@
where $\mathbf{v}$ is the whitened residual, so that the inner product of time domain residuals can be written
@@
Whitening time-domain signals via the matrix operation in (8) is generally extremely costly. However, due to being able to downsample our data to a high degree, a fairly efficient algorithm is possible for computing only the required template samples (to within some chosen degree of accuracy regarding the number of ‘significant’ neighbouring samples, see Section 3.1.4 below) to produce the downsampled residual
@@
where $\mathbf{k}$ is a sample selection vector. Then, for example, if it can be shown that the approximation method where we find the optimal noise reduction factor $m$ given in (7) is accurate, then we can write the approximate inner product as
@@ 3.1.4 Maximum correlated samples
The (approximate number of) maximum correlated samples (MCS), denoted $\mathcal{M}$, is the smallest index value of the cumulative sum of the absolute value of the (one-sided) whitening function (CSAWF), that is greater than 97% of the CSAWF at infinity.
For any given signal & detector model, this definition ensures that at least 97% of the strain value of a given whitened residual datapoint is properly accounted for. Some such cutoff is required: without it, one might need to compute all datapoints for the pre-whitening, rendering downsampling unpragmatic. For realistic detector noise profiles giving autocorrelation functions that decay smoothly, the remaining few percent are very likely to contribute close to zero to the strain value via the decorrelation.
Any given signal has its own Nyquist rate. This determines the frequency bin size/sampling time interval, and so the resolution of the ACF, thus the MCS is generally different for each signal. If some system has an MCS of 2, then 5 elements of $\mathbf{r}$ will be required to be known for each ‘1’ in $\mathbf{k}$: one corresponding to the ‘1’ in $\mathbf{k}$, the two preceding samples and the two succeeding samples. For example, if $k^{76}=1$, then we are required to know $r^{74},...,r^{78}$ in order to accurately compute $v_{N_{\mathrm{f}}}^{76}$. In general, $12\mathcal{M}+$ 1 elements of $\mathbf{r}$ must be computed for each ‘1’ in $\mathbf{k}$. This is approximate however: suppose also that $k^{77}=1$, we then need $r^{75},...,r^{79}$, but $r^{75},...,r^{78}$ will already have been found from the requirements of having $k^{76}=1$. We shall denote the approximate number of samples evaluated by $N_{\mathcal{M}}$, which, for $N_{\mathrm{s}}\ll N_{\mathrm{f}}$, is given by
@@
Note however that as $N_{\mathrm{s}}\to N_{\mathrm{f}}$, then $N_{\mathcal{M}}(N_{\mathrm{s}})\to N_{\mathrm{f}}$.
@@ 3.2 Likelihood approximation methods
To reproduce the likelihood function via downsampling, it is necessary to find methods to approximate the likelihood function. The approach we pursue is to essentially try to reshape the initial guess of the downsampled likelihood into the same form as the fully sampled one, by correcting, or compensating, for the change that occurs from discarding samples. We attempt the correction by matching the standardised moments (mean, variance, skewness, and so on) in part at least, to those of the original likelihood.
For simplicity, let us consider posterior defined by data consisting of a signal and a vanishing noise realisation. The first moment of the distribution is the location of the maximum likelihood value, which is automatically conserved under downsampling; each element of the residual vanishes at the true parameters for vanishing noise. The second moment of the distribution is then what we shall consider as the first-order contribution to the variation of the posterior from downsampling. This is known as the parameter covariance matrix (PCM). Its inverse is well-known as the Fisher information matrix (FIM), given by [17, 18]:
@@ 3.3 A single noise reduction factor
We conjecture that a single noise reweighting factor applied to all remaining samples may be sufficient to recover the general structure of the distribution after downsampling, given enough samples remain to capture a fair representation of the information content.
To introduce the motivation of this conjecture, consider the simple model, $\mathbf{h}(\theta)\!=\!\text{const}$. Suppose the detector produces time series $\mathbf{d}\!=\!\mathbf{s}+\mathbf{n}$, where $\mathbf{n}$ is a realisation of the detector noise, the signal $\mathbf{s}=\mathbf{h}(\hat{\theta})$, and $\hat{\theta}$ is the true signal parameter. Assuming Gaussian noise, the log likelihood function (discarding a constant that is of no consequence) of an original (‘full’) data set can be written as
[…]
where $\mathbf{r}_{\mathrm{f}}=\mathbf{r}_{\mathrm{f}}(\bm{\theta})\equiv\mathbf{d}_{\mathrm{f}}-\mathbf{h}_{\mathrm{f}}(\bm{\theta})$ and $N_{\mathrm{f}}=\dim(\mathbf{d}_{\mathrm{f}})$. The noise covariance matrix here is denoted $\mathbb{C}_{\mathrm{f}}$. For white noise, $\mathbb{C}_{\mathrm{f}}\!=\!\mathds{1}\sigma_{\mathrm{f}}^{2}$, where $\sigma_{\mathrm{f}}$ is the noise standard deviation. The data vector $\mathbf{d}_{\mathrm{f}}\!=\!\mathbf{s}_{\mathrm{f}}+\mathbf{n}_{\mathrm{f}}$, and suppose that $\mathbf{s}_{\mathrm{f}}\!=\!\mathbf{1}_{\mathrm{f}}\cdot\hat{\theta}$, where $\hat{\theta}$ is a constant. The model
[…]
then the log-likelihood function can be written as
$\ell_{\mathrm{f}}(\theta)$ $=-\frac{1}{2\sigma_{\mathrm{f}}^{2}}\left(\hat{\theta}^{2}\mathbf{1}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{1}_{\mathrm{f}}+\mathbf{n}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{n}_{\mathrm{f}}+2\hat{\theta}\mathbf{1}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{n}_{\mathrm{f}}\right.$
$\hskip 45.0pt-\left.2\hat{\theta}\mathbf{n}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{h}_{\mathrm{f}}-2\hat{\theta}\mathbf{1}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{h}_{\mathrm{f}}+\mathbf{h}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{h}_{\mathrm{f}}\right)$
$\approx-\frac{N_{\mathrm{f}}}{2\sigma_{\mathrm{f}}^{2}}\theta\left(\theta-2\hat{\theta}\right)\,,$
where we have discarded another constant and assumed that $\langle\mathbf{n}\rangle\approx 0$. Following the same procedure for a downsampled dataset, we find that
$\ell_{\mathrm{s}}(\theta)$ $\approx-\frac{N_{\mathrm{s}}}{2\sigma_{\mathrm{s}}^{2}}\theta\left(\theta-2\hat{\theta}\right)\,,$
and thus, as we initially set out to do, we can set $\ell_{\mathrm{f}}(\theta)\approx\ell_{\mathrm{s}}(\theta)$ if we choose:
@@
to be the noise standard deviation for the downsampled dataset.
We can try to extend this sort of analysis to the signals and detector noise profile we are interested in, however, the complexity increases very quickly for realistic physical systems, and closed form modifications become unfeasible. It is also particularly useful to have a more general procedure that is model and noise profile independent, to widen the scope of applicability.
Two (model independent) noise reduction factors are derived in the following subsections, in which we approximate posteriors as Gaussian distributions, minimise the variation of (some of) their general properties, and apply that same modification to the downsampled posterior. In the high signal to noise ratio (SNR) limit, where posteriors closely approximate Gaussians, this will be a good approximation (if the semi-axes of the Gaussians are well-aligned and proportionately similar). In other words, it is expected that only the overall precision of the likelihood will be significantly affected by downsampling. We suppose that, for our restricted class of signals of interest (LISA inspiral only signals) this assumption will generally be accurate. Since the parameter covariance matrices (PCM s) can be used to describe the aspect ratio of a $k$-dimensional ellipsoid representing the 1-sigma contour of the probability distribution [9], one might already guess that if the two ellipsoids are similar in orientation and aspect ratio, then the noise reduction factor required is the one that scales the ellipsoids to be equal in size or volume. This is of course correct; for completeness we give details in Section 3.3.1. Then for the more robust treatment in Section 3.3.2, we will also compute the factor which accounts for the Gaussians being misaligned with respect to each other and/or having unequal aspect ratios.
@@ 3.3.1 PCM volume method
Here we assume the semi-axes of the downsampled PCM s/posterior have the same orientations and aspect ratios as the PCM of the full dataset. The FIM s, $\mathbf{F}^{\prime}_{\mathrm{s}}$ $\mathbf{F}_{\mathrm{f}}$, respectively, would thus be similar, too, differing only by a scaling factor. That is, ${m}_{\mathrm{det}}\mathbf{F}^{\prime}_{\mathrm{s}}=\mathbf{F}_{\mathrm{f}}$, and therefore:
@@
where $k$ is the number of parameters. For some parameters, very little Fisher information exists, raising numerical issues; these must be projected out of the parameter space before computing $m_{\mathrm{det}}$.
@@ 3.3.2 General Gaussian approximations
Without assuming the downsampled FIM is aligned and proportionately similar to the full dataset FIM, we derive a noise reduction factor for the subspace inner product. The approach to doing this is to minimise the so-called Jeffreys’ divergence between posteriors approximated by concentric, but otherwise general $k$-dimensional Gaussians; see Appendix B for derivation and details. The noise reduction factor, from equation (89), is:
@@
where $\mathbf{D}_{\mathrm{s}}$ diagonalises $\mathbf{F}^{\prime}_{\mathrm{s}}$ and $\mathbf{D}_{\mathrm{f}}$ diagonalises $\mathbf{F}_{\mathrm{f}}$. Again, some parameters with very little information must be projected out to prevent numerical issues. With preliminary tests we found that $m_{\mathrm{J}}\sim m_{\mathrm{det}}$ when the number of datapoints used was not very low; we shall use $m_{\mathrm{J}}$ as it is an optimisation deriving from well-understood concepts of distance between distributions.
@@ 3.4 Preserving Fisher information
Given an $N_{\mathrm{f}}$-dimensional data space and metric pair $(\mathcal{D}_{\mathrm{f}},\mathbf{C}^{-1}_{\mathrm{f}})$ with FIM $F_{\mathrm{f},ij}$, we suppose there exists an $N_{\mathrm{s}}$-dimensional data space and metric pair $(\mathcal{D}_{\mathrm{s}},\mathbf{C}^{-1}_{\mathrm{s}})$ with FIM $F_{\mathrm{s},ij}$, such that
$F_{\mathrm{s},ij}=F_{\mathrm{f},ij}\,.$ (20)
Then we have
${F}_{\mathrm{s},ij}$ $={h}_{k,i}C^{-1}_{\mathrm{f},kl}{h}_{l,j}$
$={h^{\prime}}_{k^{\prime},i}C^{-1}_{\mathrm{s},kl}{h^{\prime}}_{l^{\prime},j}$
where primes indicate objects on the $N_{\mathrm{s}}$-dimensional data subspace, the $\bar{h^{\prime}}_{k^{\prime},i}$ are the prewhitened model derivative vectors on $\mathcal{D}_{\mathrm{s}}$, and $\Gamma^{-1}_{\mathrm{s}}$ is the (to-be-determined) subspace metric in the whitened basis. Given a downsampling matrix $\mathbf{D}$, for example, as in equation (4), we could write
where $\omega$ is a vector to be determined, and $\mathbf{Q}$ is some guess of sample covariance matrix. Since the samples are prewhitened, we may choose $\mathbf{Q}$ to be diagonal, and take as an ansatz $\mathbf{Q}=\mathds{1}$. With this ansatz, we can set the FIM s equal, however, there may be some other $\mathbf{Q}$ that minimises the variation between the original and downsampled posterior over the entire parameter space. Proceeding with $\mathbf{Q}=\mathds{1}$, we have that
=\;= $\sum_{k^{\prime}=0}^{N_{\mathrm{s}}}\omega_{k^{\prime}}^{2}\bar{h}^{\prime}_{k^{\prime},i}\bar{h}^{\prime}_{k^{\prime},j}$
=\;= $(\bar{h}^{\prime})^{\mathrm{\scriptscriptstyle T}}_{,i}\cdot\mathrm{diag}(\omega_{1}^{2},\omega_{2}^{2},...,\omega^{2}_{N_{\mathrm{s}}})\cdot(\bar{h}^{\prime})_{,j}\,,$ (24)
which we can compare with equation (21) to see that
$\omega^{2}_{k^{\prime}}=\sum_{p=0}^{n_{b}-1}a_{p}t^{p}_{k^{\prime}}\,,$ (26)
This can be written in matrix form $\mathbf{f}=\mathbf{Ba}$:
$\ell=-\tfrac{1}{2}(\bar{\mathbf{r}}\odot\mathbf{k})^{\mathrm{\scriptscriptstyle T}}\bm{\Gamma}^{-1}_{\mathrm{s}}(\bar{\mathbf{r}}\odot\mathbf{k})\,,$ (29)
where again $\bar{\mathbf{r}}$ is the whitened residual and $\mathbf{k}$ is the sample selection vector.
The single noise reduction factor approach discussed in the previous section is equivalent to setting the FIM s approximately equal, with the assumption that $a_{p}\!\approx\!0$ for $p\!>\!0$ and finding the $a_{0}$ that minimises the variation in the FIM s. If a situation occurs in which this single factor is not precise enough to recover lost information, which could be expected when the Fisher information transmission rate changes significantly on timescales similar to the average duration between time samples [^1] In other words, the signal cannot be said to be slowly evolving. We can define Fisher information transmission rate as the time derivative of the expectation value of the Fisher information: $\partial_{t}\langle F_{ij}(t)\rangle$., it is possible that the solution described here will be more robust (an interesting case to consider, for example, is extreme mass ratio inspirals in LISA; the rate of transmission of Fisher information is far less uniform than slow evolution BHB s, and approximate FIM recovery using a single factor may not be feasible).

### v2

@@ 3.1.2 Residuals and metrics
There are a number of ways to express the downsampled residual in terms of the fully sampled residual. First, consider the $N_{\mathrm{s}}\times N_{\mathrm{f}}$ matrix
@@
for all $\mathbf{v}$, where $\bm{\Gamma}^{-1}_{\mathrm{s}}$ and $\bm{\Gamma}^{-1}_{\mathrm{f}}$ are inner products on the subspace and full data space respectively. In general, an exact solution for $\bm{\Gamma}^{-1}_{\mathrm{s}}$ would depend on the full residual vector $\mathbf{v}$. Our method will be to make an initial guess of a metric on the subspace ${\bm{\Gamma}^{\prime}}_{\mathrm{s}}^{-1}$, and define the final metric $\bm{\Gamma}^{-1}_{\mathrm{s}}=\bm{\Gamma}^{-1}_{\mathrm{s}}(\bm{\Gamma}^{\prime}_{\mathrm{s}})$ in terms of the initial guess in some appropriate fashion, as shall be described in the following sections. With ${\bm{\Gamma}^{\prime}}_{\mathrm{s}}^{-1}$ an $N_{\mathrm{s}}\times N_{\mathrm{s}}$ matrix, we can project the full data space metric down to the subspace using $\mathbf{D}$ via
@@
3.1.3 Time-domain sample correlations
@@
The above example shows how correlations cause the information in the time-domain signal to which the detector is sensitive to become mixed between neighbouring samples. In order to prevent this from impinging upon the Bayesian inference after downsampling, we must whiten, or in some other way decorrelate the data samples (for example, frequency representations from fast Fourier transform (FFT) s often consist of uncorrelated samples). Thus in fact the preceding prescription for downsampling given in Section 3.1.2 should only be carried out on uncorrelated data: since frequency-domain data is naturally uncorrelated, the task would be straightforward and safe (at least from the point of view of sample correlations) to downsample frequency samples. In the time-domain, we must first whiten residuals before downsampling. This simply means setting
@@
where $\mathbf{r}$ is the time-domain residual and $\mathbf{v}$ its whitened counterpart, so that the inner product of time-domain residuals can be written
@@
Whitening time-domain signals via the matrix operation in (8) is generally extremely costly. However, due to being able to downsample our data to a high degree, an efficient algorithm is possible for computing only the required template samples (to within some chosen degree of accuracy regarding the number of ‘significant’ neighbouring samples, see Section 3.1.4 below) to produce the downsampled residual
@@
where $\mathbf{k}$ is a sample selection vector. If it can be shown that the approximation method where we find the optimal noise reduction factor $m$ given in (7) is accurate, for example, then we can write the approximate inner product as
@@ 3.1.4 Maximum correlated samples
The (approximate number of) maximum correlated samples (MCS), denoted $\mathcal{M}$, is the smallest index of the cumulative sum of the absolute value of the (one-sided) whitening function (CSAWF) for which the cumulative sum exceeds 97% of its asymptotic value.
For any given signal–detector model, this definition ensures that at least 97% of the contribution to a given whitened residual sample is explicitly accounted for. Some cutoff of this kind is essential: without it, one would in principle need to compute all data points to perform the pre-whitening, which would defeat the purpose of downsampling. For realistic detector noise profiles, whose autocorrelation functions decay smoothly, the remaining few percent of the whitening weight is concentrated at large lags and is therefore expected to contribute negligibly to the residual value.
Any given signal has its own Nyquist rate. This fixes the sampling interval (and hence the frequency-bin size), which in turn determines the resolution of the ACF; the MCS is therefore generally different for each signal. If some system has an MCS of $\mathcal{M}=2$, then $2\mathcal{M}+1=5$ elements of $\mathbf{r}$ must be known for each ‘1’ in $\mathbf{k}$: the sample corresponding to the ‘1’ and its two immediate predecessors and successors. For example, if $k^{76}=1$, then we require $r^{74},\dots,r^{78}$ in order to compute $v^{76}$ accurately. In general, $2\mathcal{M}+1$ elements of $\mathbf{r}$ must be computed for each ‘1’ in $\mathbf{k}$.
This counting is only approximate, however, because the required neighbourhoods can overlap. For instance, if also $k^{77}=1$, then we nominally need $r^{75},\dots,r^{79}$, but $r^{75},\dots,r^{78}$ have already been obtained from the requirements of $k^{76}=1$. We denote the approximate total number of template samples that must be evaluated by $N_{\mathcal{M}}$, which for $N_{\mathrm{s}}\ll N_{\mathrm{f}}$ is well approximated by
@@ 3.2 Likelihood approximation methods
To reproduce the effect of the full-data likelihood under downsampling, we require a way to approximate the likelihood evaluated on the reduced dataset. Our approach is to start from a natural downsampled likelihood and then ‘reshape’ it towards the full-data likelihood by correcting for the change induced by discarding samples. We attempt this correction by matching, at least in part, the second-order structure of the posterior. In the Gaussian/high-SNR regime, this is encoded by the parameter covariance matrix (or equivalently the Fisher information matrix). In one approach we adjust a single noise-reweighting factor so that a scalar summary of this structure (the overall volume of the covariance ellipsoid) is preserved under downsampling. In a more general approach we take the full Fisher matrices into account.
For intuition, let us first consider the posterior defined by data consisting of a signal plus a zero-noise realisation. In this case the residual $\mathbf{r}(\bm{\theta})$ vanishes at the true parameters, so the maximum-likelihood point is unchanged by downsampling: modifying the inner product or discarding samples does not move the point where $\mathbf{r}=0$. The leading effect of downsampling is therefore on the curvature of the log likelihood around this maximum, i.e. on the second-order structure of the posterior.
In the Gaussian/high-SNR regime this second-order structure is encoded in the parameter covariance matrix (PCM), whose entries are the second central moments of the posterior. The PCM is well approximated by the inverse of the Fisher information matrix (FIM), given by [FisherDef, FisherDef2]:
@@ 3.3 A single noise reduction factor
We conjecture that a single noise reweighting factor applied to all remaining samples may be sufficient to recover the general structure of the distribution after downsampling, provided enough samples remain to capture a fair representation of the information content.
To motivate this conjecture, consider the simple model $\mathbf{h}(\theta)=\theta\mathbf{1}_{\mathrm{f}}$, where $\mathbf{1}_{\mathrm{f}}$ is an $N_{\mathrm{f}}$-dimensional vector of ones. Suppose the detector produces time series $\mathbf{d}_{\mathrm{f}}=\mathbf{s}_{\mathrm{f}}+\mathbf{n}_{\mathrm{f}}$, with $\mathbf{s}_{\mathrm{f}}=\mathbf{h}(\hat{\theta})=\hat{\theta}\,\mathbf{1}_{\mathrm{f}}$, where $\hat{\theta}$ is the true parameter, and $\mathbf{n}_{\mathrm{f}}$ is a realisation of white Gaussian noise with covariance $\mathbb{C}_{\mathrm{f}}=\sigma_{\mathrm{f}}^{2}\mathds{1}$. The residual is
$\mathbf{r}_{\mathrm{f}}(\theta)=\mathbf{d}_{\mathrm{f}}-\mathbf{h}_{\mathrm{f}}(\theta)=\mathbf{n}_{\mathrm{f}}+(\hat{\theta}-\theta)\mathbf{1}_{\mathrm{f}}\,.$
Ignoring additive constants, the log likelihood is
$\ell_{\mathrm{f}}(\theta)\propto-\frac{1}{2\sigma_{\mathrm{f}}^{2}}\,\mathbf{r}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}(\theta)\,\mathbf{r}_{\mathrm{f}}(\theta)\,.$
Expanding the norm gives
$\mathbf{r}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{r}_{\mathrm{f}}=\mathbf{n}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{n}_{\mathrm{f}}+(\hat{\theta}-\theta)^{2}\mathbf{1}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{1}_{\mathrm{f}}+2(\hat{\theta}-\theta)\,\mathbf{1}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{n}_{\mathrm{f}}\,.$
For zero-noise data or upon averaging over noise realisations, the cross term vanishes and $\mathbf{1}_{\mathrm{f}}^{\mathrm{\scriptscriptstyle T}}\mathbf{1}_{\mathrm{f}}=N_{\mathrm{f}}$, so we obtain
$\ell_{\mathrm{f}}(\theta)\approx-\frac{N_{\mathrm{f}}}{2\sigma_{\mathrm{f}}^{2}}(\theta-\hat{\theta})^{2}\,.$
Performing the same construction for a downsampled dataset of size $N_{\mathrm{s}}$ with white noise variance $\sigma_{\mathrm{s}}^{2}$, we find
$\ell_{\mathrm{s}}(\theta)\approx-\frac{N_{\mathrm{s}}}{2\sigma_{\mathrm{s}}^{2}}(\theta-\hat{\theta})^{2}\,.$
To match the curvature of the two log likelihoods we can choose
@@
so that $\ell_{\mathrm{s}}(\theta)\approx\ell_{\mathrm{f}}(\theta)$.
We can try to extend this sort of analysis to the signals and detector noise profiles of interest. However, the complexity increases very quickly for realistic physical systems, and closed-form modifications become unfeasible. It is also desirable to have a more general procedure that is model- and noise-profile independent.
Two (model-independent) noise reduction factors are derived in the following subsections, in which we approximate posteriors as Gaussian distributions, minimise the variation of certain scalar summaries of their second-order structure, and apply the corresponding modification to the downsampled posterior. In the high-signal to noise ratio (SNR) limit, where posteriors closely approximate Gaussians, this should be a good approximation (provided the semi-axes of the Gaussians are reasonably well aligned and proportionately similar). In other words, we expect that, for our restricted class of signals of interest (LISA inspiral-only signals), downsampling mainly affects the overall precision of the likelihood rather than its shape.
Since a parameter covariance matrix (PCM) describes the aspect ratio of the $k$-dimensional ellipsoid associated with the 1-sigma contours of the distributions [UseAbuseFisher], one might already guess that if the two ellipsoids are similar in orientation and aspect ratio, then the appropriate noise reduction factor is the one that rescales their volumes to agree. This is indeed the case; details are given in Section 3.3.1. For a more robust treatment in Section 3.3.2, we also compute a factor that accounts for the Gaussians being misaligned and/or having unequal aspect ratios.
@@ 3.3.1 PCM volume method
Here we assume that the semi-axes of the downsampled PCM (and hence of the downsampled posterior) have the same orientations and aspect ratios as those of the full-data PCM. The corresponding FIM s, $\mathbf{F}^{\prime}_{\mathrm{s}}$ and $\mathbf{F}_{\mathrm{f}}$, then differ only by an overall scaling factor, so that
$m_{\mathrm{det}}\,\mathbf{F}^{\prime}_{\mathrm{s}}=\mathbf{F}_{\mathrm{f}}\,.$
Taking determinants gives
@@
where $k$ is the number of parameters. For some parameters, very little Fisher information exists, which can lead to numerical issues; these directions must be projected out of the parameter space before computing $m_{\mathrm{det}}$.
@@ 3.3.2 General Gaussian approximations
Without assuming that the downsampled FIM is aligned and proportionately similar to the full-data FIM, we instead derive a noise reduction factor for the subspace inner product by minimising the so-called Jeffreys’ divergence between posteriors approximated by concentric but otherwise general $k$-dimensional Gaussians; see Appendix A for the derivation and details. The resulting noise reduction factor, from equation (59), is
@@
where $\mathbf{D}_{\mathrm{s}}$ diagonalises $\mathbf{F}^{\prime}_{\mathrm{s}}$ and $\mathbf{D}_{\mathrm{f}}$ diagonalises $\mathbf{F}_{\mathrm{f}}$. Again, parameters with very little Fisher information must be projected out to prevent numerical issues. In preliminary tests we found that $m_{\mathrm{J}}\sim m_{\mathrm{det}}$ whenever the number of retained data points was not extremely small; we shall nevertheless use $m_{\mathrm{J}}$, since it arises from an optimisation based on a standard notion of distance between distributions.
@@ 3.4 Preserving Fisher information
There exist Fisher-information-lossless encodings of the data under downsampling, as we show in Appendix B; that is, the FIM can be preserved exactly after downsampling. In this construction the downsampled log likelihood is written as in equation (69), reproduced here:
$\ell=-\tfrac{1}{2}(\mathbf{D}\bar{\mathbf{r}})^{\mathrm{\scriptscriptstyle T}}\bm{\Gamma}^{-1}_{\mathrm{s}}(\mathbf{D}\bar{\mathbf{r}})\,$,
where $\bar{\mathbf{r}}$ is the whitened residual, $\mathbf{D}$ is the sample selection matrix, and $\bm{\Gamma}^{-1}_{\mathrm{s}}$ is the inner-product matrix on the whitened, downsampled signal space, as derived in Appendix B.